40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 217 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  6426   Fri Mar 16 16:03:03 2012 kiwamuUpdateIOOMC alignment servo : put some offsets in the TRANS QPD signal

The MC alignment servo wasn't great in the last 1 hour or so as it kept disturbing the MC lock. It was found to be due to some offsets in the MC trans QPD signals.

I put some values to cancel the offsets and then the lock became stable.

This is a first aid. So we need to take a closer look at the QPD signals and also probably the spot position on the QPD.

 


The symptom was that every time the alignment servo was engaged, at the beginning the amount of the transmitted light went to 27000 counts, which is good.

However, then the amount of the transmitted light slowly decreased in a time scale of ~ 20 sec or so, ending up with destruction of the MC lock.

According to the time scale I suspected that the servos using the trans QPD signals were doing something bad because their control width had been designed to be slow and slower than the rest of the servo loops.

I switched off the servos, called C1:IOO-TRANS_PIT and C1:IOO-TRANS_YAW and found the MC stayed locked stably with 27000 counts of the transmitted light.

Leaving the trans QPD servos off, I zeroed the offsets and then switched them on. It worked.

 

The values below are the current offset that I put.

                C1:IOO-MC2_TRANS_PIT_OFFSET = -0.115203
                C1:IOO-MC2_TRANS_YAW_OFFSET = -0.0323576
 

  6425   Fri Mar 16 16:01:53 2012 ranaUpdateElectronicsREFL165 calibration : measurements

 To characterize the RF V to counts we need to know the state of the whitening filter board. Was the filter on or off ? What was the value of the whitening gain slider?

  6424   Fri Mar 16 10:37:52 2012 JenneUpdateElectronicsJenne Laser

Quote:

Here is a picture of the small black breadboard on which I have put together the PD characterisation setup.  It would be great if we can retain this portable set up as it is, since we keep reusing it every couple of weeks.  It would be convenient if we can fiber couple the path to the PD under test with a 2m long fiber.  Then we will not have to remove the PD from the optical table while testing it. 

 This is totally sweet Suresh!  I don't remember how much more fiber is coiled up under the plate that has the "Jenne Laser" label, but there's a reasonable amount.  It's not 2m, but maybe we can just extend the blue snakey thing some?

  6423   Fri Mar 16 06:17:56 2012 SureshUpdateElectronicsREFL165 calibration : measurements

 

These are the measurements for estimating the amplitude of the signal recorded in the CDS when a known amount of modulated light is incident on the photodiode. 

I mounted the PD characterisation setup onto a small breadboard which could then be placed close AP table.  I then placed position markers for REFL165 on the AP table before moving it onto my small breadboard.  The AM laser was driven by an RF function generator (Fluke 6061A) at a frequency of 165.98866 MHz, which is 102 Hz offset from the 165MHz LO.  The power level was set at -45dBm.  This power level was chosen since anything higher would have saturated the AntiAliasing  Whitening Filters.  The counts in the CDS were converted to voltage using the ADC resolution = 20V per 2^16 counts.

  

  RF source RF power to AM laser 1611 PD 1611 PD REFL165 REFL165 CDS CDS
  power set (dBm) Actual power out (dBm) DC (V) RF out (dBm) DC (mV) RF out (dBm) Amplitude (V)   102 Hz Amplitude (V) 102 Hz
                 
1  -45  -50.6  -2.5 -58.9  10  -37.4  0.171 0.172
2  -48  -53.5  -2.5 -62.1  10  -40.3  0.122  0.121
3  -51  -56.5  -2.5 -65.0  10  -43.1  0.085  0.085

    

 When the 166MHz power is decreased by a factor of 2 the amplitude of 102Hz wave recorded in CDS goes down by sqrt(2) as expected.   The RF AM power incident on the REFL165 was estimated to be 0.011mW(rms)  (case #1 in the above table)  using the DC power ratio and using the transimpedance of the 1611 BBPD to be 700 Ohms.  This produces a 171 mV amplitude wave at 102 Hz.  I then stepped down the power by factor of 2 and repeated the measurement. 

(These numbers however are not agreeing with the power incident on REFL165 if we assume its transimpedance to be 12500.  It will take a bit more effort to make all the numbers agree.  Will try again tomorrow)

Here is a picture of the small black breadboard on which I have put together the PD characterisation setup.  It would be great if we can retain this portable set up as it is, since we keep reusing it every couple of weeks.  It would be convenient if we can fiber couple the path to the PD under test with a 2m long fiber.  Then we will not have to remove the PD from the optical table while testing it.

IMG_0552.JPG

 

  6422   Thu Mar 15 08:48:40 2012 RyanSummaryCDSSummary of Syracuse Visit to 40m Mar 5-9 2012

JIMS Channels in PEM Model

The PEM model has been modified now to include the JIMS(Joint Information Management System) channel processing. Additionally Jim added test points at the outputs of the BLRMS.

For each seismometer channel, five bands are compared to threshold values to produce boolean results.  Bands with RMS below threshold produce bits with value 1, above threshold results in 0.  These bits are combined to produce one output channel that contains all of the results.

A persistent version of the channel is generated by a new library block that called persist which holds the value at 0 for a number of time steps equal to an EPICS variable setting from the time the boolean first drops to zero. The persist allows excursions shorter than the timestep of a downsampled timeseries to be seen reliably.

The EPICS variables for the thresholds are of the form (in order of increasing frequency):
C1:PEM-JIMS_GUR1X_THRES1
C1:PEM-JIMS_GUR1X_THRES2
etc.
 
The EPICS variables for the persist step size are of the form:
C1:PEM-JIMS_GUR1X_PERSIST
C1:PEM-JIMS_GUR1Y_PERSIST
etc.

The JIMS Channels are being recorded and written to frames:

The two JIMS channels at 2048:
[C1:PEM-JIMS_CH1_DQ] Persistent version of JIMS channel. When bit drops to zero indicating something bad (BLRMS threshold exceeded) happens the bit stays at zero  for >= the value of the persist EPICS variable.
[C1:PEM-JIMS_CH2_DQ] Non-persistent version of JIMS channel.

And all of the BLRMS channels at 256:
Names are of the form:
[C1:PEM-RMS_ACC1_F0p1_0p3_DQ]
[C1:PEM-RMS_ACC1_F0p3_1_DQ]

For additional details about the JIMS Channels and the implementation, please see the previous elog entries by Jim.

 

Conlog

I have a working aLIGO Conlog/EPICS Log installed and running on megatron.

Please see this wiki page for the details of use:
https://wiki-40m.ligo.caltech.edu/aLIGO%20EPICs%20log%20%28conlog%29

I also edited this page with restart instructions for megatron:
https://wiki-40m.ligo.caltech.edu/Computer_Restart_Procedures#megatron

Please see Ryan's previous elog entries for installation details.

Future Work

  • Determine useful thresholds for each band
  • Generate MEDM Screens for JIMS Channels
  • Add a decimation option to channels
  • Add EPICS Strings in PEM model to describe bits in JIMS Channels
  • Add additional JIMS Channels: Testing additional characterization methods
  • Implement a State Log on Megatron: Will Provide a 1Hz index into JIMS Channels
  • Generate a single web page that allows access to aLIGO Conlog/EPICS Log and State Log

 

  6421   Thu Mar 15 04:04:23 2012 KojiConfigurationLockingPRC Matching issues

Kiwamu and Koji

We found that the intra-cavity mode of the PRC is not round although it was obvious even with the DARK and REFL port images.
We need to review the mode matching situation.

In order to look at the PRC intra-cavity mode, we reconfigured the POP CCD.

If we look at the beam reflected from the Michelson, the beam is round. However, the PRC intra-cavity mode can never be round
in any resonant conditions. (Pict 1, 2, and 3, for the sideband resonant, carrier resonant conditions and another carrier resonant
one, respactively). Particularly the mode of the carrier resonant case is very unstable and always changing.

P3150902.jpg

P3150906.jpgP3150907.jpg

By misaligning the PRM, we can compare between the spot directly reflected from the Michelson and the one after additional round trip in the PRC (Pic 4).

They looks round, but it was obvious the secondary reflection is dimmer and larger (Pic 5). The intensity difference corresponds to the factor RPRM RMI
(i.e. product of the reflectivities for the PRM and MI). It can be understand if the dimmer spot looks smaller due to the artifact of the CCD. But it is opposite.

This may mean the mode matching is not correct. We are not sure what is not right. This could be just an incorrect incident beam, the curvature error of the PRM,
beam is distortec by the TT mirrors, or some other unknown reasons.

More precise analysis can be done with quantitative analysis of those two spots with Beamscan. This could happen tomorrow.

P3150910.jpgP3150900.jpg

  6420   Wed Mar 14 23:02:09 2012 KojiUpdateLSCLocking activity

Kiwamu and Koji

The target is to realize DRMI or PRMI + one arm with ALS.

The focus of the night is to achive stable lock of the PRMI (SB resonant) with 3f signals.
Particularly, REFL165 is back now, we are aiming to see if any of the 165 signals is useful.

We made a comparison between  REFL33Q/REFL165Q/AS55Q to find any good source of MICH.
However, none of them showed a reasonable shape of the spectra. They don't have reasonable coherence between them.

Nonetheless, we have tried to lock the IFO with those REFL signals. But any of them were useful to keep the PRMI (SB resonant).
The only kind of stable signal for MICH was AS55Q as we could keep the PRMI locked.

  6419   Wed Mar 14 21:01:36 2012 keikoUpdateLSCevolution of the sensing matrix in PRMI as a function of time

This is the simulated signals to compare with the original post #6403

 

 

PRMI configuration, PRCL signal

[W/m] Simulation Measured
REFL11 575440

 

~10000

REFL33 4571 ~50
REFL55 288400 ~5000
REFL165 891 NA
AS55 71 70

 

PRMI configuration, MICH signal

[W/m] Simulation Measured
REFL11 2290

 

~600

REFL33 36 ~4
REFL55 5623 ~200
REFL165 17 NA
AS55 6456 ~200
 

Simulated DC REFL power is 9mW (before the attenuator). AS DC is 0.3mW.

They don't agree. I suspect the PR gain for the SBs are somehow different. It is about 40 (or a bit less) in the simulation for 11MHz.

 

 

 

 

  6418   Wed Mar 14 16:39:02 2012 SureshUpdateGeneralREFL165 signal was not reaching demod board : Fixed

Quote:

The following tasks need to be done in the daytime tomorrow.

  • Hook up the DC output of the Y green BBPD on the PSL table to an ADC channel (Jamie / Steve)
  • Install fancy suspension matrices on PRM and ITMX [#6365] (Jenne)
  • Check if the REFL165 RFPD is healthy or not (Suresh / Koji)
    • According to a simulation the REFL165 demod signal should show similar amount of the signal to that of REFL33.
    • But right now it is showing super tiny signals [#6403]

 The REFL165 RF output was not reaching the Demod board.  The RF cable was disconnected.  I fixed that and then I put in a RF signal at 165MHz , 1.66 mVrms at the test input  (100Hz off set from the 165MHz LO) and saw that the 100 Hz demodulated signal was visible in the dataviewer. 

Test_CDS_Calibration.png

 

Will complete the Optical RF power -> CDS counts calibration tomorrow morning. 

  6417   Wed Mar 14 16:33:20 2012 keikoUpdateLSCRAM simulation / RAM pollution plot

In the last post, I showed that SRCL element in the MICH sensor (AS55I-mich) is chaned 1% due to RAM.

Here I calculated how is this 1% residual in MICH sensor (AS55 I-mich) shown in MICH sensitivity. The senario is:

(1) we assume we are canceling SRCL in MICH by feed forward first (original matrix (2,3) element).

(2) SRCL in MICH (matrix(2,3) is changed 1% due to RAM, but you keep the same feed forward with the same feedforward gain

(3) You get 1% SRCL residual motion in MICH sensor. This motion depends on how SRCL is quiet/loud. The assumed level is

Pollution level = SRCL shot noise level in SRCL sensor  x  SRCL closed loop TF  x  1% residual .... the following plot.

 

 

AS sensor = AS55I-mich  --- SN level 2.4e-11 W/rtHz ------- MICH SN level 6e-17 m/rtHz

SRCL sensor = AS55 I-SRCL --- SN level 2e-11 W/rtHz ---  SRCL SN level 5e-14 m/rtHz

 

 

RAMexampleplot.png

 

 

Quote:

Adding some more results with more realistic RAM level assumption.

(4) With 0.1% RAM mod index of PM (normalized by (1) )

  PRCL MICH SRCL
REFL11I 0.99999 -0.001807 -0.000148
AS 55 Im 0.000822 1.000002 0.000475
AS 55 Is 1.068342 906.968167 1.00559
 

 

 

 

 

Attachment 1: Mar14pollution.png
Mar14pollution.png
  6416   Wed Mar 14 14:09:01 2012 interferomter tack forceUpdateGeneraldaytime tasks

Quote:

The following tasks need to be done in the daytime tomorrow.

  • Hook up the DC output of the Y green BBPD on the PSL table to an ADC channel (Jamie / Steve)
  • Install fancy suspension matrices on PRM and ITMX [#6365] (Jenne)
  • Check if the REFL165 RFPD is healthy or not (Suresh / Koji)
    • According to a simulation the REFL165 demod signal should show similar amount of the signal to that of REFL33.
    • But right now it is showing super tiny signals [#6403]

 For ITMX, I used the values from the conlog:

2011/08/12,20:10:12 utc 'C1:SUS[-_]ITMX[-_]INMATRIX'
These are the latest values in the conlog that aren't the basic matricies.  Even though we did a round of diagonalization in Sept, and the 
matricies are saved in a .mat file, it doesn't look like we used the ITMX matrix from that time.

For PRM, I used the matricies that were saved in InputMatricies_16Sept2011.mat, in the peakFit folder, since I couldn't find anything in the Conlog other than the basic matricies.

 

UPDATE:  I didn't actually count the number of oscillations until the optics were damped, so I don't have an actual number for the Q, but I feel good about the damping, after having kicked POS of both ITMX and PRM and watching the sensors.

  6415   Wed Mar 14 13:27:15 2012 ZachHowToComputer Scripts / ProgramsModeMatchr

I have added to ModeMatchr the capability to fix the total MMT distance. This is nice if you are coupling to a cavity some fixed distance away. The blurb from the help:

% Note: for any total length constraint dtot_tol > 0, ModeMatchr will use
% fminsearch to find the best solutions near your nominal dtot, and then
% omit solutions whose dtot lie outside your tolerance. For dtot_tol = 0,
% ModeMatchr actively constrains dtot to your value, and then finds the
% best solution. Therefore, set dtot_tol = 0 if you have a fixed distance
% into which to put a MMT.

Quote:

For your mode matching pleasure, I have added a tool called "ModeMatchr" to the SVN under /trunk/zach/tools/modematchr/

It uses the usual fminsearch approach, but tolerates a fully astigmatic input (i.e., w0ix ≠ w0iy, z0ix ≠ z0iy) and allows for transforming to an elliptical waist  (i.e., w0fx ≠ w0fy, but z0fx = z0fy). It would be straightforward to allow for z0fx ≠ z0fy, but I have never seen a case when we actually wanted this. On the other hand, the elliptical output ability is nice for coupling to wide-angle ring cavities.

It also does the looping through available lenses for you , and retains the best solution for each lens combination in an output cell, which can then be combed with another function (getOtherSol). fminsearch is incredibly fast: with a 10-lens bank, it finds all 100 best solutions on my crappy MacBook in <10s.

I have also included the functionality to constrain the length of the total MMT to within some percentage of the optimal distance, which helps to sift through the muck .

MMrLogo.png

 

  6414   Wed Mar 14 13:16:50 2012 kiwamuUpdateLSCA correction on Noise estimatino in the REFL33Q

A correction on the previous elog about the REFL33Q noise:

 Rana pointed out that the whitening filter's input referred noise should not be such high (I have estimated it to be at 54 nV/sqrtHz).
In fact the measurement was done in a condition where no laser is on the photo diode by closing the mechanical shutter at the PSL table.
Therefore the noise I called "whitening filter input referred noise" includes the voltage noise from the RFPD and it could have such a noise level.
So the noise curve drawn in the plot should be called "whitening filter + RFPD electronics noise".

Quote from #6407

A feasibility study of the REFL33Q as a MICH sensor was coarsely performed from the point view of the noise performance.

  • Whitening filter input referred noise
    • I assumed that it is flat with a level of 54 nV/sqrtHz based on a rough measurement by looking at the spectrum of the LSC input signals.
    • The contribution was estimated by applying some gain corrections from the conversion efficiency of the demod board, transimpedance gain, responsivity and the optical gain.
    • This noise is currently the limiting factor over a frequency range from DC to 1 kHz.

 

  6413   Wed Mar 14 10:06:26 2012 steveUpdateGreen Lockingdichroic mirror quotes

Dichroic mirror quotes are in the wiki.

ATF is pricy.

We got a good price from Laseroptik, but the wedges are 5 arcminutes. The fused silica grade is 0F, meaning the  homogeneity is 5 ppm instead of 1ppm.  I requested an other large wedge quote on the substrates.We may have to get substrates from somebody else and ship it to Germany

MLT quote is outrageously high

REO is not interested in this low volume job.

 

  6412   Wed Mar 14 05:26:39 2012 interferomter tack forceUpdateGeneraldaytime tasks

The following tasks need to be done in the daytime tomorrow.

  • Hook up the DC output of the Y green BBPD on the PSL table to an ADC channel (Jamie / Steve)
  • Install fancy suspension matrices on PRM and ITMX [#6365] (Jenne)
  • Check if the REFL165 RFPD is healthy or not (Suresh / Koji)
    • According to a simulation the REFL165 demod signal should show similar amount of the signal to that of REFL33.
    • But right now it is showing super tiny signals [#6403]
  6411   Wed Mar 14 04:19:51 2012 kiwamuUpdateLSCREFL33Q for MICH control : not good

 I tried the REFL33Q for controlling MICH in the PRMI configuration (#6407)

The result was --

 It was barely able to lock MICH in a short moment but didn't stay locked for more than 10 sec. Not good.

 

The attached screenshot below shows a moment when the PRMI was locked with REFL33I and REFL33Q for PRCL and MICH respectively.
Apparently the lock was destroyed after 10 sec or so and it was locked again.
Untitled.png

 

(Tricks)

 At the beginning I tried minimizing the PRCL signal in the Q phase by rotating the demodulation phase because the PRCL signal was always huge.
However it turned out that the rotation of the demodulation phase didn't completely eliminate the PRCL signal for some reason.
 
This could be some kind of imbalance in the electronics or somewhere between the I and Q signal paths.
So instead, I tried blending the I and Q signals by a linear combination through the LSC input matrix.
Then I was able to eliminate the PRCL signal.
I put a gain of -0.1 for the I signal and 1 for the Q signal to get the good blend when the demodulation phase was at -17.05 deg.
  6410   Wed Mar 14 04:03:37 2012 kiwamuUpdateIOOPZT1 and associate extra works

As the PZT1 has not been functional, I have been aligning the Y arm to the input beam instead of aligning the beam to the Y arm.

It turned out that this procedure leads to two extra works everytime after alignments of the Y arm:

  1. The Y green beam must be always aligned to the Y arm
    • The amount of the misalignment was found to be relatively big compared with how it used to be.
  2. The PSL beat note setup must be always realigned because the Y green path is determined by the orientation of the Y arm.
    • In the past I didn't often realign the beat note path, but currently it needs to be pay more attentions.

Sad ..

Quote from #6357

   The polarity for controlling the PZT1 PITCH seems to have flipped for some reason.

 

  6409   Wed Mar 14 03:34:44 2012 kiwamuUpdateSUSAdjustment of BS suspension output matrix : coupling from SIDE to POS

[Rana / Kiwamu]

 We put some elements in the BS output matrix to mitigate the actuator coupling from SIDE to POS.

As a result the degree of the coupling reduced by a factor of 2 or so.

Rana did the "Q of 5" test on the SIDE damping servo after putting the elements and set the gain to be 40.

 

The attached screen shot is the new elements that we put in the suspension output matrix.

Untitled.png

 

(How to)

  • Excite the SIDE motion by AWG at 3 Hz.
  • Monitor the POS signal in DTT
  • Try some numbers in the matrix elements until the peak at 3 Hz in the POS signal is minimized

Quote from #6369

The BS SIDE damping gain seemed too low. The gain had been 5 while the rest of the suspensions had gains of 90-500.

I increased the gain and set it to be 80.

 

I did the "Q of 5" test by kicking the BS SIDE motion to find the right gain value.

However there was a big cross coupling, which was most likely a coupling from the SIDE actuator to the POS motion.

Due to the cross coupling, the Q of 5 test didn't really show a nice ring down time series. I just put a gain of 80 to let the Q value sort of 5.

I think we should diagonalize the out matrices for all the suspensions at some point.

 

  6407   Tue Mar 13 19:14:40 2012 kiwamuUpdateLSCNoise estimatino in the REFL33Q as a MICH sensor

A feasibility study of the REFL33Q as a MICH sensor was coarsely performed from the point view of the noise performance.

The answer is that :

  the REFL33Q can be BARELY used as a MICH sensor in the PRMI configuration, but the noise level will be at only sub-nano meter level.

  Tonight I will try to use the REFL33Q to control the MICH DOF to see what happens.

 

(Background)

  I neeeeeeeed a 3f signal which is sensitive enough to hold the Michelson in the PRMI configuration so that I can test the single arm + PRMI configuration.
Based on the data I got in the sensing matrix measurement (#6403) I wanted to see how noises in the REFL33Q look like.
 

(Noise analysis)

  I did a coarse noise analysis for the REFL33Q signal as shown in the attached plot below while making some assumptions as follows.

  •  Optical gain for MICH = 0.8  W/m (#6403)
    • In the plot below, I plotted a unsuppressed MICH motion which had been measured the other day with a different sensor. This is for a comparison.
  •  Shot noise due to DC light on the REFL33 photo diode
    •  With a power of 5.0 mW (#6355)
    • Assume that the responsivity is 0.75 A/W, this DC light creates the shot noise in the photo current at a level of 35 pA/sqrtHz.
    • Then I estimated the contribution of this shot noise in terms of the MICH displacement by calibrating the number with the optical gain and responsivity.
    • It is estimated to be at 60 pm/sqrtHz
  • Dark current
    • I assumed that the dark current is 0.52 mA. (see the wiki)
    • In the same manner as that for the shot noise, the dark current is estimated to be at 20 pm/sqrtHz in terms of the displacement
  • Whitening filter input referred noise
    • I assumed that it is flat with a level of 54 nV/sqrtHz based on a rough measurement by looking at the spectrum of the LSC input signals.
    • The contribution was estimated by applying some gain corrections from the conversion efficiency of the demod board, transimpedance gain, responsivity and the optical gain.
    • This noise is currently the limiting factor over a frequency range from DC to 1 kHz.
  • ADC noise
    • I did the same thing as that for the whitening filter noise.
    • I assumed the noise level is at 6 uV/sqrtHz and it is flat (I know this not true particularly at mHz region the noise becomes bigger by some factors)
    • Then I applied the transfer function of the whitening filter to roll off the noise above 15 Hz.

 NB_REFL33.png

(Some thoughts)

  •   Obviously the limiting noises are that of ADC and the whitening filter.
    • These noise can be easily mitigated by installing an RF amplifier to amplify the RF signals from the REFL33Q RFPD.
    • Therefore this is not the real issue
  • The real issue is that the shot noise is already at a level of 60 pm/sqrtHz, and we can't suppress the MICH motion less than that.
    • In order to decrease it, one possibility is to increase the modulation depth. But it is already at the maximum.
    • If the REFL165 RFPD is healthy, it is supposed to give us a bigger MICH signal. But it didn't look healthy ... (#6403)
  6406   Tue Mar 13 16:56:19 2012 kiwamuUpdateLSCevolution of the sensing matrix in PRMI as a function of time

Next steps:

  • Compare the obtained sensing matrix with an Optickle model. Particularly I am interested in the absolute strengths in watts/meter
  • Noise estimation of the REFL33_Q as a MICH sensor to see if this sensor is usable for holding MICH.

Quote from #6403

Tonight I measured the sensing matrix again but this time I recorded them as a function of time using the realtime LOCKINs in the LSC front end.

The attached plots are the responses (optical gains) of PRCL and MICH in watts / meter at various sensors in time series.

  6405   Tue Mar 13 16:40:06 2012 kiwamuUpdateLSCevolution of the sensing matrix in PRMI as a function of time: details

Here I describe the measurement of the sensing matrix.

 

Motivations

  There were two reasons why I have been measuring the sensing matrix :

  1.  I wanted to know how much each element in the sensing matrix drifted as a function of time because the sensing matrix didn't agree with what Optickle predicted (#6283).
  2.  I needed to estimate the MICH responses in the 3f demodulated signals, so that I can decide which 3f signal I should use for holding MICH.

 I will report #2 later because it needs another careful noise estimation.

 

Measurement

 In order to measure the sensing matrix, the basic steps are something like this:

  1. Excite one of the DOF at a certain frequency, where a notch filter is applied in the LSC servos so that the servos won't suppress the excitation signal.
  2. Demodulate the LSC signals (e.g. C1:LSC-REFL11_I_ERR and etc.,) by the realtime LOCKINs (#6152) at the same frequency.
  3. Calibrate the obtained LOCKIN outputs to watts/meter.
In the actual measurement I choose the frequency of the excitation signal to be at 283.1 Hz,
at which any of the LSC servos don't have gains of more than 1 and there were no particular structures in the spectra.
For the amplitude of the excitation, I usually choose it to be 1000 - 2000 counts.
Because all the actuators have response functions of approximately 10-9 / f^2 meter/counts  (#5637), the actual displacement in the excited DOF should be about 10 pm level.
Therefore the excited displacements must be always in the linear ranges and also the amplitude in counts is reasonably smaller than the DAC range.
 

LOCKIN detection

The attached cartoon below shows how the LOCKIN system works for the MICH response measurement.
In the case of the PRCL response measurement, the setup is the same except that only PRM is shaken.
Here is some notes about the LOCKIN detection.
  • The LOCKIN oscillator excites ITMs differentially
    • In order to purely excites the MICH DOF, the actuation coefficients were precisely adjusted (#6398).
    • Currently ITMY has a gain of 1, and ITMX has a gain of -0.992 for the pure MICH excitation. Those numbers were put in the output matrix of the LOCKIN oscillator.
  • The demodulation phase of the LOCKIN system was adjusted to be -22 deg at the digital phase rotator.
    • This number maximizes the in-phase signals while the quadrature-phase signals give almost zero.
    • This number was adjusted when the simple MICH configuration was applied.
  • In the demodulations, the LO signals have amplitude of 100 counts to just make the demodulated signals readable numbers.

 

lockins_MICH.png

 

Calibration of the LOCKINs

  The calibration of the LOCKIN detectors is easy because all the processes takes place in the digital land, where we know all the parameters.
In this phase the goal is to calibrate the signals into counts / meter.
To calibrate the LOCKIN output signals, the following equation is used :
 
 [The obtained LOCKIN output in counts ] = H x ADOF x CLO x CEXC x 1/2  ,
 
 where H is the response of a sensor (e.g. AS55_I, AS55_Q and so on) against a particular DOF in unit of counts / m and this the quantity which we want to measure here,
ADOF is the actuator efficiency of the DOF at the excitation frequency in unit of m/counts,
CLO is the amplitude of the local oscillator signal for demodulating the sensor signals in unit of counts,
CEXC is the amplitude of the excitation signal in unit of counts,
the last 1/2 term comes from the fact there is a low pass filter in each demodulation path. 
Therefore once we measure the response of a sensor, dividing the obtained LOCKIN output by ADOF x CLO x CEXC x 1/2 gives the calibrated response in unit of counts/meter.
  ADOF are well known as they have been measured several times (#5637).
For the MICH actuator I assumed that AMICH = 2 x (ITMY response) since they are balanced through the actuation coefficients.
Note that a confirmation of this calibration has been done
when the configuration is in the simple Michelson, where we can easily estimate the response of a sensor by letting the MICH freely swing.
 

Calibration of the responses to watts/meter

  With the calibration process described above, we obtain the sensor responses in unit of counts/m.
 Then we need to do another calibration to make them into unit of W/m.
If we think about how the RFPD signal flows, we get the following gain chain.
 
[raw response in counts/m ] = Hopt x CADC x Ldemod x GWF x Ztrans x RPD
 
Hopt  is the optical gain at a sensor which we want to calibrate. It is in unit of W/m.
CADC  is the conversion factor of the ADCs and the value is CADC = 1638.4 counts/m because their resolution is 16 bit and the range is +/-20 V.
Ldemod is the conversion efficiency of the demodulation boards in unit of V/V. I used the values which Suresh measured yesterday (#6402).
GWF is the gain of the whitening filter in unit of V/V,
Ztrans is the transimpedance gain of an RFPD in unit of V/A and I used the values summarized in (the wiki),
and RPD is the responsivity of the photo diodes and I assumed RPD = 0.75 A/W for all the RFPDs.
 
Therefore the calibration can be done by dividing the raw response value by the entire gain chain of CADC x Ldemod x GWF x Ztrans x RPD.
 

Settings and parameters

  •  LSC RF demodulation phases
    •  AS55 = 17.05 deg (minimizing the PRCL sensitivity in the Q-phase)
    •  REFL11 = -41.05 deg (maximizing the PRCL sensitivity in the I-phase)
    • REFL33 = -25.85 deg (maximizing the PRCL sensitivity in the I-phase)
    • REFL55 = 4 deg (maximizing the PRCL sensitivity in the I-phase)
    • REFL165 = 39 deg (random number)
  •  Whitening filters
    • AS55 = 30 dB
    • REFL11 = 0 dB
    • REFL33 = 42 dB
    • REFL55 = 30 dB
    • REFL165 = 45 dB
  • MICH servo
    • AS55_Q for the sensor
    • G = -5 in the digital gain
    • FM2, FM3, FM5 and FM9 actiavted
    • UGF ~ 100 Hz
    • Feedback to ITMs differentially
  • PRCL servo
    • REFL33_I for the sensor
    • G = 1 in the digital gain
    • FM2, FM3, FM4, FM5 and FM9 activated
    • UGF ~ 100 Hz
    • Feedback to PRM

Quote from #6403

Tonight I measured the sensing matrix again but this time I recorded them as a function of time using the realtime LOCKINs in the LSC front end.

I will explain some more details about how I measured and calibrated the data in another elog entry.

  6404   Tue Mar 13 13:28:31 2012 Ryan FisherUpdateCDSDAQ restart with new ini file
Extra note: This was the ini file that was edited:
/cvs/cds/rtcds/caltech/c1/chans/daq/C1PEM.ini
  6403   Tue Mar 13 07:04:55 2012 kiwamuUpdateLSCevolution of the sensing matrix in PRMI as a function of time

The punch line is -- the sensing matrix still looks strange in the PRMI configuration.

 

I have been measuring the sensing matrix of the PRMI configuration because it didn't make sense (#6283).

One strange thing I have noticed before was that all the I-phase signals showed a weird behavior -- they fluctuate too much in time series.

Tonight I measured the sensing matrix again but this time I recorded them as a function of time using the realtime LOCKINs in the LSC front end.

The attached plots are the responses (optical gains) of PRCL and MICH in watts / meter at various sensors in time series.

I will explain some more details about how I measured and calibrated the data in another elog entry.

 

PRCL.png

 MICH.png

 

  6402   Mon Mar 12 22:14:56 2012 SureshUpdateRF SystemCalibration of Demod Board Efficiency.

I have completed the calibration of the demod board efficiencies.  Here is the schematic of the set-up.

 Calibration_Schematic.png

The data is given below and the data-file is attached in several different formats.

 Demod_calib.png

 

Attachment 3: Demod_calib.txt
								
	Measurements			 After corrections			Efficiency= out/in	
Demod Board	mV_ampl	mV_pk-pk	mV_pk-pk	mV_ampl	mV_ampl	mV_ampl	Vout/Vin	Vout/Vin
	PD in	Q out	I out	PD in	Q out	I out	Q out	I out
REFL33	10.6	10.0	10.0	9.4	5.0	5.0	0.53	0.53
AS11	24.0	10.0	11.0	21.3	5.0	5.5	0.23	0.26
REFL11	22.5	240.0	255.0	20.0	120.0	127.5	6.00	6.38
POX11	24.0	9.2	8.5	21.3	4.6	4.3	0.22	0.20
POY11	22.4	10.5	9.0	19.9	5.3	4.5	0.26	0.23
AS55	17.6	268.0	268.0	15.6	134.0	134.0	8.57	8.57
REFL55	19.7	15.8	15.5	17.5	7.9	7.8	0.45	0.44
POP55	18.8	278.0	274.0	16.7	139.0	137.0	8.32	8.20
REFL165	21.2	16.0	16.4	18.8	8.0	8.2	0.42	0.44
POY110	23.4	14.7	14.4	20.8	7.4	7.2	0.35	0.35
POY22	17.5	11.9	9.3	15.6	6.0	4.6	0.38	0.30
Attachment 4: Demod_calib.xlsx
  6401   Mon Mar 12 18:57:58 2012 keikoUpdateLSCRAM simulation update, RAM LSC matrix

Quote:

 I calculated the DRMI RAM LSC matrix with RAM and the operation point offsets.

  • configuration: C1 DRMI
  • RAM is added by an Mach-Zehnder ifo placed before the PRM
  • demodulation phases are optimised for each DoF
  • the operation points offset from the PDH signals are calculated and added to the optical configuration as mirror position offsets
  • Then the matrix is calculated with the offsets and the RAM
  • The set of the scrips are found as RAMmatrix.m, normMAT.m, newGetMAT.m,  on CVS/ifomodeling/40m/fullIFO_Optickle. They are a bit messy scripts at this moment.

Results:

(1) No RAM LSC matrix

  PRCL MICH SRCL
REFL11I 1 -0.001806 -0.000147
AS 55Q 0.000818 1 0.000474
AS 55 I 1.064561 902.292816 1

(2) With 1% RAM mod index of PM (normalised by (1) )

  PRCL MICH SRCL
REFL11I 1.000618 -0.001837 -0.000163
AS 55Q 0.000919 1.000521 0.000495
AS 55 I 1.169741 924.675187 1.018479
 

(3) With 5% RAM mod index of PM (normalised by (1) )

  PRCL MICH SRCL
REFL11I 0.999986 -0.001812 -0.000150
AS 55Q 0.000838 1.000028 0.000479
AS 55 I 1.084598 906.83668 1.003759
 

Adding some more results with more realistic RAM level assumption.

(4) With 0.1% RAM mod index of PM (normalized by (1) )

  PRCL MICH SRCL
REFL11I 0.99999 -0.001807 -0.000148
AS 55Q 0.000822 1.000002 0.000475
AS 55 I 1.068342 906.968167 1.00559
 

(5) With 0.5% RAM mod index of  PM (normalized by (1) )

  PRCL MICH SRCL
REFL11I  0.999978  -0.001810    -0.000149 
AS 55Q 0.000830  1.000010  0.000476 
AS 55 I 1.075926 904.321433  1.001677
 

  6400   Mon Mar 12 01:04:18 2012 keikoUpdateLSCRAM simulation update, RAM LSC matrix

 I calculated the DRMI RAM LSC matrix with RAM and the operation point offsets.

  • configuration: C1 DRMI
  • RAM is added by an Mach-Zehnder ifo placed before the PRM
  • demodulation phases are optimised for each DoF
  • the operation points offset from the PDH signals are calculated and added to the optical configuration as mirror position offsets
  • Then the matrix is calculated with the offsets and the RAM
  • The set of the scrips are found as RAMmatrix.m, normMAT.m, newGetMAT.m,  on CVS/ifomodeling/40m/fullIFO_Optickle. They are a bit messy scripts at this moment.

Results:

(1) No RAM LSC matrix

  PRCL MICH SRCL
REFL11I 1 -0.001806 -0.000147
AS 55Q 0.000818 1 0.000474
AS 55 I 1.064561 902.292816 1

(2) With 1% RAM mod index of PM (normalised by (1) )

  PRCL MICH SRCL
REFL11I 1.000618 -0.001837 -0.000163
AS 55Q 0.000919 1.000521 0.000495
AS 55 I 1.169741 924.675187 1.018479
 

(3) With 5% RAM mod index of PM (normalised by (1) )

  PRCL MICH SRCL
REFL11I 0.999986 -0.001812 -0.000150
AS 55Q 0.000838 1.000028 0.000479
AS 55 I 1.084598 906.83668 1.003759
 

  6399   Sat Mar 10 15:29:47 2012 ZachHowToComputer Scripts / ProgramsModeMatchr

For your mode matching pleasure, I have added a tool called "ModeMatchr" to the SVN under /trunk/zach/tools/modematchr/

It uses the usual fminsearch approach, but tolerates a fully astigmatic input (i.e., w0ix ≠ w0iy, z0ix ≠ z0iy) and allows for transforming to an elliptical waist  (i.e., w0fx ≠ w0fy, but z0fx = z0fy). It would be straightforward to allow for z0fx ≠ z0fy, but I have never seen a case when we actually wanted this. On the other hand, the elliptical output ability is nice for coupling to wide-angle ring cavities.

It also does the looping through available lenses for you , and retains the best solution for each lens combination in an output cell, which can then be combed with another function (getOtherSol). fminsearch is incredibly fast: with a 10-lens bank, it finds all 100 best solutions on my crappy MacBook in <10s.

I have also included the functionality to constrain the length of the total MMT to within some percentage of the optimal distance, which helps to sift through the muck .

MMrLogo.png

  6398   Sat Mar 10 02:00:03 2012 keikoUpdateLSCupdate on the locking activity

ITMX and ITMY balance for the MICH excitation (lockin) is adjusted again. Now it's ITMx = -0.992, ITMy = 1 for MICH (lockin output matrix values).

RA: what were the old values? Does this change make any difference for the signal mixing noticed before?

  6397   Fri Mar 9 20:44:24 2012 Jim LoughUpdateCDSDAQ restart with new ini file

DAQ reload/restart was performed at about 1315 PST today. The previous ini file was backed up as c1pem20120309.ini in the /chans/daq/working_backups/ directory.

I set the following to record:

The two JIMS channels at 2048:
[C1:PEM-JIMS_CH1_DQ] Persistent version of JIMS channel. When bit drops to zero indicating something bad (BLRMS threshold exceeded) happens the bit stays at zero  for >= the value of the persist EPICS variable.
[C1:PEM-JIMS_CH2_DQ] Non-persistent version of JIMS channel.

And all of the BLRMS channels at 256:
Names are of the form:
[C1:PEM-RMS_ACC1_F0p1_0p3_DQ]
[C1:PEM-RMS_ACC1_F0p3_1_DQ]

On monday I intend to look at the weekend seismic data to establish thresholds on the JIMS channels.

256 was the lowest rate possible according to the RCG manual. The JIMS channels are recorded at 2048 because I couldn't figure out how to disable the decimation filter. I will look into this further.

  6396   Fri Mar 9 16:28:10 2012 Ryan FisherSummaryComputer Scripts / ProgramsAlterations to base epics install for installing aLIGO conlog:
I created a page on the wiki for the new EPICS log (conlog):
https://wiki-40m.ligo.caltech.edu/aLIGO%20EPICs%20log%20%28conlog%29

I also edited this with restart instructions:
https://wiki-40m.ligo.caltech.edu/Computer_Restart_Procedures#megatron
  6395   Fri Mar 9 16:00:46 2012 steveUpdateGreen Lockinglaser emergency shut down switch replaced at the south end

Over-sized local laser emergency switch was held by large C clamp at the south end. This was replaced by a smaller one and it is mounted with magnets.

The Innolight laser was turned off, while the interlock was wired.

  6394   Fri Mar 9 15:48:56 2012 Ryan FisherSummaryComputer Scripts / ProgramsAlterations to base epics install for installing aLIGO conlog:

I decided to make a backup of the database and then delete it and make a new database:
 
cd ~/ryan/database_dumpMar92012
mysqldump -u root -p C1_conlog > C1_conlog.dump.Mar92012 
Note: it appears this failed the first time, thankfully this wasn't a production service yet... In the future, do not trust this backup method for important data!

Next, log into mysql as root, dump the database, remake it and grant privileges again.:
(This is saved in megatron:~/ryan/restore_database.txt
megatron:~/ryan>mysql -u root -p
Enter password:
Welcome to the MySQL monitor.  Commands end with ; or \g.
Your MySQL connection id is 174
Server version: 5.1.41-3ubuntu12.10 (Ubuntu)

Type 'help;' or '\h' for help. Type '\c' to clear the current input statement.

mysql> list databases;
ERROR 1064 (42000): You have an error in your SQL syntax; check the manual that corresponds to your MySQL server version for the right syntax to use near 'list databases' at line 1
mysql> list users;      ERROR 1064 (42000): You have an error in your SQL syntax; check the manual that corresponds to your MySQL server version for the right syntax to use near 'list users' at line 1
mysql> use C1_conlog
Reading table information for completion of table and column names
You can turn off this feature to get a quicker startup with -A

Database changed
mysql> list users;
ERROR 1064 (42000): You have an error in your SQL syntax; check the manual that corresponds to your MySQL server version for the right syntax to use near 'list users' at line 1
mysql> select User from mysql.user;                                             +------------------+
| User             |
+------------------+
| php              |
| C1_conlog_epics  |
| c1_conlog_epics  |
| root             |
| C1_conlog_epics  |
| c1_conlog_epics  |
| debian-sys-maint |
| root             |
| root             |
+------------------+
9 rows in set (0.00 sec)

mysql> show databases;                                                          +--------------------+
| Database           |
+--------------------+
| information_schema |
| C1_conlog          |
| mysql              |
+--------------------+
3 rows in set (0.00 sec)

mysql> drop database C1_conlog ;
Query OK, 2 rows affected (0.56 sec)

mysql> create database C1_conlog;
Query OK, 1 row affected (0.00 sec)

mysql> use C1_conlog ;
Database changed
mysql> SET SQL_MODE="NO_AUTO_VALUE_ON_ZERO";
Query OK, 0 rows affected (0.00 sec)

mysql>
mysql> CREATE TABLE `channels` (
    ->   `channel_id` mediumint(8) unsigned NOT NULL AUTO_INCREMENT,
    ->   `channel_name` varchar(60) NOT NULL,
    ->   PRIMARY KEY (`channel_id`),
    ->   UNIQUE KEY `channel_name` (`channel_name`)
    -> ) ENGINE=MyISAM  DEFAULT CHARSET=latin1;
Query OK, 0 rows affected (0.04 sec)

mysql>
mysql> CREATE TABLE `data` (
    ->   `acquire_time` decimal(26,6) NOT NULL,
    ->   `channel_id` mediumint(8) unsigned NOT NULL,
    ->   `value` varchar(40) DEFAULT NULL,
    ->   `status` tinyint(3) unsigned DEFAULT NULL,
    ->   `connected` tinyint(1) unsigned NOT NULL,
    ->   PRIMARY KEY (`channel_id`,`acquire_time`)
    -> ) ENGINE=MyISAM DEFAULT CHARSET=latin1;
Query OK, 0 rows affected (0.03 sec)

mysql> grant select, insert, update, execute on * to 'c1_conlog_epics'@'127.0.0.1';  Query OK, 0 rows affected (0.00 sec) 
mysql> grant select, insert, update, execute on * to 'C1_conlog_epics'@'127.0.0.1';   Query OK, 0 rows affected (0.00 sec) 
 mysql> grant select, insert, update, execute on * to 'c1_conlog_epics'@'localhost';  Query OK, 0 rows affected (0.00 sec) 
mysql> grant select, insert, update, execute on * to 'C1_conlog_epics'@'localhost';
Query OK, 0 rows affected (0.00 sec)

mysql> grant select on C1_conlog to 'php'@'%';
ERROR 1146 (42S02): Table 'C1_conlog.C1_conlog' doesn't exist
mysql> grant select on * to 'php'@'%';
Query OK, 0 rows affected (0.00 sec)

mysql> select * from mysql.users
    -> ;
ERROR 1146 (42S02): Table 'mysql.users' doesn't exist
mysql> select User from mysql.user;        
| C1_conlog_epics  |
| c1_conlog_epics  |
| root             |
| C1_conlog_epics  |
| c1_conlog_epics  |
| debian-sys-maint |
| root             |
| root             |
+------------------+
9 rows in set (0.00 sec)

mysql> Bye 



Next, I decided that I want to index on the acquire_time instead of the combination of channel_id and acquire_time (I think it makes a lot of sense for several query types, and especially debugging the conlog!):
mysql> create index acquire_time_index on data(acquire_time);
Query OK, 0 rows affected (0.04 sec)
Records: 0  Duplicates: 0  Warnings: 0

Next Fix:


The above worked well, but when I restarted the conlog, I had to re-execute the "remove_channels" from the medm, because initially all channels were being loaded (use_channel_names had all the channels still).
Additionally, there were a lot of channels with "*RMS*" in the name that were being recorded, and were changing relatively quickly, so I have added those to the remove_channel_names file.

I am going to: Backup the files in /ligo/caltech/data/conlog/c1
Edit use_channel_names to only have the good channels.
Dump the database again
Stop conlog.
Wipe the database again.
Remake the database again (with permissions and the new index).
Restart the conlog and hope!

The fix above seems to be in place and working. The database has the initial entries for the channels it monitors and is not growing without operators changing EPICs values.

  6393   Fri Mar 9 13:34:13 2012 keikoUpdateLSCupdate on the locking activity

We tried to measure the sensing matrix for MICH and PRCL last night. They look too much mixed as we expect... the matrix may be posted later. We suspect the IX and IY of the MICH excitation is not balanced very well, although Kiwamu adjusted that about two weeks ago, and it is mixing the dof. We'll try to balance it again, ans see the matrix. 

Keiko, Kiwamu

 

Quote:

[Keiko / Kiwamu]

 Some updates on the locking activity:

  • Started summarizing the data of the Michelson lock in a wiki page:
  • Gradually moving on to the PRMI lock
    • The lock stays for reasonably a long time (~20 min or more)
    • POP22/110 demod signals seemed just ADC noise.
    • A first noise budget is in process
      • The glitches make the noise level worse above 40 Hz or so in both the MICH and PRCL budgets.
    • Sensing matrix will be measured tomorrow
    • The data will be also summarized in a wiki page

 

  6392   Fri Mar 9 11:59:38 2012 Zweizig the ELOG MavenSummaryCDSNDS2 restart

 Hi Rana,

It looks like the channe list file has a few blank lines that the channel list reader is choking on. I removed the lines and it is working now.. I have made the error message a bit more obvious (gave the file name  and line number) and allowed it to ignore empty lines so this won't cause problems with future versions (when installed). The bottom line is nds2 is now running on mafalda.
Best regards,

John   

  6391   Fri Mar 9 11:02:56 2012 Ryan FisherSummaryComputer Scripts / ProgramsAlterations to base epics install for installing aLIGO conlog:

Too Many Fast Changing EPICS in New Conlog


I have been monitoring the new conlog, and it already has far too many rows.

I'm going through the list of channels to exclude in the update_channels script for the conlog that is currently running and removing them from
the monitored channels in the new conlog using the remove_channel_names file and the medm screen (we may want to just wipe out the tables
and start over after this is set properly, but for now I'm keeping them):
#-- Exclude a few uninteresting or obsolete categories
if ( $chan =~ m/^[BIJ]$/ ||
$chan =~ m/IOO-MC_PWR_IN/ ||
$chan eq "C1:PSL-FSS_SLOWDC" ||
$chan =~ m/PSL-STAT_.*_BITS/ ||
$chan =~ m/:IOO-PZTM[12]_(PIT|YAW)_BIAS$/ ||
$chan =~ m/DAQ.*_cycle/ ||
$chan =~ m/DAQ.*rtSeconds/ ||
$chan =~ m/C1:-.*/ ||
$chan =~ m/C1:SUP/ ||
$chan =~ m/C1:SP/ ||
$chan =~ m/C1:X/ ||
$chan =~ m/C1:TST/ ||
$chan =~ m/C1:RF/ ||
$chan =~ m/C1:UCT/ ||
$chan =~ m/C1:DU/ ||
$chan =~ m/C1:MCP/ ||
$chan =~ m/C1:MCS/ ||
$chan =~ m/C1:FEC/ ||
$chan =~ m/C1:PEM/ ||
$chan =~ m/C1:LSP/ ||
$chan =~ m/C1:NIO/ ||
$chan =~ m/C1:WFS/ ||
$chan =~ m/C1:ASC-WFS/ ||
$chan =~ m/C1:ASC-SP/ ||
$chan =~ m/C1:VG/ ||
$chan =~ m/C1:IOO-DOF/ ||
$chan =~ m/C1:IOO-EO/ ||
$chan =~ m/Name/ ||
$chan =~ m/DEFAULTNAME/ ||
$chan =~ m/:IOO-PZT.*OFFSET/ ||
$chan =~ m/PD_VAR$/ ||
$chan =~ m/_INMON$/ ||
$chan =~ m/_EXCMON$/ ||
$chan =~ m/_OUT16$/ ||
$chan =~ m/_OUTMON$/ ||
$chan =~ m/_OUTPUT$/ ||
$chan =~ m/_RSET$/ ||
$chan =~ m/_ALIVE$/ ||
$chan =~ m/VMon$/ ||
$chan =~ m/PDMon$/ ||
$chan =~ m/(BiasVMon|FE_PPOLL|MASTER_OVERFLOW|FSS_TIDALSET|CPU_LOAD|CDM
_STAT|State_Bits|INDCOFFSET)/ )

With these removals, only 15493 channels are being monitored now.
  6390   Fri Mar 9 10:44:57 2012 steveUpdateRF SystemOSA

Optical spectrum analyzers like the Attachment made by Coherent , Meles Griot- CVI and Spectral Product are all discontinued.

The 40m have Coherent models C240 analyzer with controller C251 Their Finesse measured in 2004: sn205408  F302,  sn205409  F396,

Jenne borrowed Jan's Meles Griot model 13SAE006, Peter King has the same model. FSR 300 MHZ, finnees 200 minimum

Attachment 1: OSA.pdf
OSA.pdf
  6389   Fri Mar 9 10:17:06 2012 steveHowToPEMdo not leave tables open

Type: How not to

Please do not leave optical tables open! You will be held responsible for creating dirty optics.

Attachment 1: P1080547.JPG
P1080547.JPG
  6388   Thu Mar 8 23:37:03 2012 kiwamuUpdateIOOdither Y arm dither script

I disabled the feedback to the PZT1 PITCH in the Y arm dithering scripts so that it won't push the beam away from the good point.

Currently one has to do a manual alignment only for the PZT PITCH but the rest of DOFs are still able to be automatically aligned with the script.

Quote from #6357

 The polarity for controlling the PZT1 PITCH seems to have flipped for some reason.

  6387   Thu Mar 8 17:18:19 2012 Ryan FisherSummaryComputer Scripts / ProgramsAlterations to base epics install for installing aLIGO conlog:
I have the aLIGO Conlog 'working' at http://192.168.113.209/conlog/conlog.php

The process is running inside a screen on megatron.

To start it running, you need to set your environment, and then run the startup.c1conlog script :

cd /cvs/cds/rtcds/caltech/c1/target/conlog/conlogepics

source conlog_environment.txt

./startup.c1conlog

This will leave you at an epics prompt, which means the code is running. (that's why I left it running in a screen for now).

To change the list of channels when the conlog is running, you need to edit the file (s):

/ligo/caltech/data/conlog/c1/add_channel_names
/ligo/caltech/data/conlog/c1/remove_channel_names

Then start up medm as follows:

cd ~/ryan/

medm -x -macro "IFO=C1" medm/CONLOG.adl

Then click the Add channel list button or Remove channel list button.

To change the channels before running the conlog from a blank database, you would edit:
/ligo/caltech/data/conlog/c1/use_channel_names (I believe this should be read whenever the conlog is restarted, but I'm only sure it does the first time you start conlog).




Documenting the rest of the installation:


Successful? Installation of Fresh EPICS and Extensions



Fresh Copy of EPICS 3.14.10


* We restarted (on Patrick's suggestion) with a fresh copy of EPICS 3.14.10 in:
/ligo/apps/linux-x86_64/epics
* I had to set a clean environment:
* Then I downloaded the tarball, unpacked it, and simply ran make within it, and it worked!
* Next, I followed Patrick's wiki instructions with only modifications to the configure/RELEASE files for the archive and ioc/conlog extensions.
* Then I realized that I had to rebuild conlog ioc after adding a directory:
/ligo/apps/linux-x86_64/epics/iocs/conlog/iocBoot/iocc1_conlog
* I had to copy the files from the h1 directory and modify them so that all reference to h1 now point to c1 in the new directory
* I then rebuilt the conlog ioc (I had to make sure setenv SCRIPTS was run again because I had logged out and back in, and I reset the whole environment properly again)

Rest of Install


* I was able to fairly trivially follow through the rest of the installation steps on Patrick's wiki, up to the "Design" section.
* Now, there is no obvious way to move forward (nothing is actually running I believe).

New Install Instructions:


"
You want to create the EPICS IOC boot directory by doing the following:

In the top level of the IOC (.../epics/iocs/conlog) with the appropriate
paths:

/epics/base/bin/linux-x86_64/makeBaseApp.pl -i -t ioc c1_conlog
What application should the IOC(s) boot?
The default uses the IOC's name, even if not listed above.
Application name? conlog

Then you will have to update the st.cmd it creates. You can compare the
st.cmd and st.cmd.backup files in the other directories to see what needs
to be changed.

I don't know if just copying the directory will work, it might.

You will also probably want to change the following line in
/epics/modules/archive/archiveApp/src/drvArchive.c:

queue = epicsMessageQueueCreate(100000000, sizeof(struct message_type));

and reduce 100000000 to something smaller depending on the amount of ram
available to the computer. I think sizeof(struct message_type) is
something like 112 bytes. Then recompile.

You basically put a file with the list of channels to use in the directory
path for "use channel list filepath" in the following command in st.cmd:
drvArchive_read_channel_list_filepaths <add channel list filepath>,<remove
channel list filepath>,<use channel list filepath>
I can send you the script that I currently use to generate that channel
list if you want, but it may need to be changed for your setup.

Once you start the ioc, open the medm which can be checked out from
subversion here: cds_user_apps/trunk/cds/common/medm/CONLOG.adl
with macro substitution for IFO: medm -x -macro "IFO=C1"
and click on the button for "Use channel list".
The number of monitored channels should increase to the number of channels
in the file.

-Patrick

...

The perl script and example file are attached, just redirect the output of
the perl script to a file. It scans autoBurt.req files in a particular
directory and its subdirectories for channel names that meet certain
criteria. All the file contains is a list of channel names, one on each
line. To start the IOC, go to the target directory and run
./startup.c1conlog.

"

{{rpfisher:scan_autoburt.pl.txt|scan_autoburt.pl.txt}}



{{rpfisher:use_channel_names.txt|use_channel_names.txt}}



My Notes Regarding These Instructions:


* Throughout the installation instructions, it probably should have been made clear that the ifo is lowercase: eg c1 (but in the end the installation mixed C1 and c1 in different places)
* Also throughout, one should be careful to replace lho with your site (eg caltech) wherever it appears
* After running the first perl script to generate the iocBoot/iocc1_conlog directory, the goal is to rebuild the whole conlog ioc by running make from epics/iocs/conlog, but before doing that:
* I needed to change the suggested line in the archive module to match the correct RAM size of the machine I am installing on (I actually gave it just less than half the free RAM), then:
* Remake the archive module
* Change into the ioc/conlog directory, remove the iocBoot I had made before for c1, rerun the perl script above, then run make from the ioc/conlog directory.
* Once that is done, you need to edit the file st.cmd to add lines for the reading and writing of channels, such as:
dbLoadRecords("db/conlog.db","IFO=C1")
drvArchive_mysql "C1_conlog","/data/mysql/C1_conlog_epics_user"
drvArchive_read_channel_list_filepaths "/ligo/caltech/data/conlog/c1/add_channel_names","/ligo/caltech/data/conlog/c1/remove_channel_names","/ligo/caltech/data/conlog/c1/use_channel_names"
drvArchive_write_channel_list_filepaths "/ligo/caltech/data/conlog/c1/channel_names","/ligo/caltech/data/conlog/c1/monitored_channel_names","/ligo/caltech/data/conlog/c1/unmonitored_channel_names"
* I also had to rerun this set of commands once that was all done:
controls: cd /opt/rtcds/<site>/<ifo>/target/conlog/conlogepics
controls: cp -r /ligo/apps/linux-x86_64/epics/iocs/conlog/db ./
controls: cp -r /ligo/apps/linux-x86_64/epics/iocs/conlog/dbd ./
controls: cp /ligo/apps/linux-x86_64/epics/iocs/conlog/iocBoot/ioc<ifo>_conlog/envPaths ./
controls: cp /ligo/apps/linux-x86_64/epics/iocs/conlog/iocBoot/ioc<ifo>_conlog/st.cmd ./
controls: echo ./bin/linux-x86_64/conlog st.cmd > startup.<ifo>conlog
controls: chmod a+x startup.<ifo>conlog
* This set of instructions seemed to be lacking, so I added this:

cp -r /ligo/apps/linux-x86_64/epics/iocs/conlog/bin/ ./

* Now the executable runs but doesn't work:
* Fixes needed:
* Need to use root permissions and make sure the files in /data/mysql have the right names for the users the code expects and also have the right passwords. (have to match capitalization appropriately for the <ifo> tag everywhere
* Might need to go into mysql and add a new user with the proper capitalization also
* Need to edit the ld_library_path to point to the new epics libraries (see the suggested environment below)
* Now, the code seems to work, but dumps me at an "epics> " prompt, I'm asking Patrick what to do next.
* I was impatient and loaded up the medm screen, and found out that the one channel I had picked was not readable (unmonitored)
* I ran a modified version of Patrick's perl script to search autoBert files for channels, and replaced my use_channel_names file with the output of the script
* Now, the medm screen shows lots of monitored channels, and the conlog is filling up! (can see it from phpmyadmin)
* Next step: I wanted to get the php pages working, so I edited the files inside /var/www/conlog:
megatron:~/ryan>diff -u /var/www/conlog/conlog.php /var/www/conlog/conlog.php.bck.Mar82012_1
--- /var/www/conlog/conlog.php  2012-03-08 15:31:53.152547771 -0800
+++ /var/www/conlog/conlog.php.bck.Mar82012_1   2012-03-08 15:28:23.062704171 -0800
@@ -19,7 +19,7 @@
        <form action="query.php" method="post">
                <h3><label for="database">Database:</label></h3>
                <select id="database" name="database">
-                       <option value="C1_conlog">C1</option>
+                       <option value="h2_conlog">h2</option>
                </select>
 
                <h3><label for="included_channels">Included channels:</label></h3>

megatron:~/ryan>diff -u /var/www/conlog/query.php /var/www/conlog/query.php.bck.Mar82012_1
--- /var/www/conlog/query.php   2012-03-08 15:33:45.122550303 -0800
+++ /var/www/conlog/query.php.bck.Mar82012_1    2012-03-08 15:32:31.772554679 -0800
@@ -168,8 +168,8 @@
        }
 
        $database_name = $_POST["database"];
-       if ($database_name == 'C1_conlog') {
-               $server = '192.168.113.209';
+       if ($database_name == 'h2_conlog') {
+               $server = 'cdsconlog';
        }
        else {
                die('Unknown database.');

* Finally, the mysql server was denying access from outside queries, so I fixed that:
megatron:~/ryan>diff -u /etc/mysql/my.cnf /etc/mysql/my.cnf.bck.Mar82012_1
--- /etc/mysql/my.cnf   2012-03-08 15:35:57.122548370 -0800
+++ /etc/mysql/my.cnf.bck.Mar82012_1    2012-03-08 15:35:10.652559315 -0800
@@ -49,7 +49,7 @@
 #
 # Instead of skip-networking the default is now to listen only on
 # localhost which is more compatible and is not less secure.
-#bind-address          = 127.0.0.1
+bind-address           = 127.0.0.1
 #
 # * Fine Tuning
 #
megatron:~/ryan>
* Now, I think everything is working * almost:
* It seems that when you first start the conlog up, it finds all the variables and inserts values of "Null" for everything, but after that it detects changes properly!


Conlog Environment


Need to source this to use the new environment:

megatron:~>cat ~/ryan/conlog_enviroment.txt
  6386   Thu Mar 8 04:13:12 2012 kiwamuUpdateLSCupdate on the locking activity

[Keiko / Kiwamu]

 Some updates on the locking activity:

  • Started summarizing the data of the Michelson lock in a wiki page:
  • Gradually moving on to the PRMI lock
    • The lock stays for reasonably a long time (~20 min or more)
    • POP22/110 demod signals seemed just ADC noise.
    • A first noise budget is in process
      • The glitches make the noise level worse above 40 Hz or so in both the MICH and PRCL budgets.
    • Sensing matrix will be measured tomorrow
    • The data will be also summarized in a wiki page
  6385   Thu Mar 8 00:57:48 2012 keikoUpdateLSCMICH noise budget on Mar 5, Mar 6, and old

Here is the recent two noise budgets of MICH, with the old measurement by Jenne. The most latest Mar 6 data is quite close to the old data, even better around 20-30 Hz. Probably some scattering source was improved?

Mar7MICHbudgettotal.png

  6384   Wed Mar 7 23:29:28 2012 keikoUpdateLSCREFL OSA observation

 kiwamu, keiko

 

 

REFLOSA.png

We measure the REFL OSA spectrum when (1) direct reflection from the PRM (2) CR lock at PRC (3) SB lock at PRC. When CR lock, both SBs are reflected from the PRC and when SB lock (ref line), some SB is sucked by PRM and looked lower than the other two lines.

 

  6383   Wed Mar 7 23:28:41 2012 Lab Cleanup CrewHowToEnvironmentTrue Beauty....

Or, how a lab should look at the end of every day.

Clean_40m_Electronics_Bench.jpg

Beat that, Bridge kids!

  6382   Wed Mar 7 22:04:05 2012 kiwamuUpdateLSCREFL OSA : how the signal look like

I was also wondering about the same thing, comparing with what Mirko obtained before with the same OSA ( #5519).

Quote from #6379

I'm puzzled why the 11MHz peak can be such high considering 1.7~2 times smaller the modulation depth.

 

  6381   Wed Mar 7 21:13:30 2012 ranaUpdateDAQNDS2

 I noticed that NDS2 was not running on mafalda as it should be. Instead, there were a couple of zombie MEDMs using up 99% of the CPU. I killed the zombies and have run the 'build channel list' script. When it finished, I tried to restart the nds server, but got the following error in the log file. Email has been dispatched to JZ.

mafalda:logs>less nds2-mafalda-201203072111.log

Configuring from file: nds2.conf
Allow list: ALL
terminate called after throwing an instance of 'std::runtime_error'
  what():  Insufficient arguments
  6380   Wed Mar 7 20:53:13 2012 keikoUpdateLSCMICH noise budget on 5 Mar

 

 Mar6-MICHbudget.png

This is the MICH noise budget on 6th March. 1Hz peak got a bit better as the BS sus control gain was increased.

 

  6379   Wed Mar 7 20:06:23 2012 KojiUpdateLSCREFL OSA : how the signal look like

I'm puzzled why the 11MHz peak can be such high considering 1.7~2 times smaller the modulation depth.

  6378   Wed Mar 7 19:10:06 2012 kiwamuUpdateLSCREFL OSA : how the signal look like

Just a quick report on the REFL OSA.

The attached plot below shows the raw signal from the REFL OSA which Keiko installed in this afternoon.

When the data was taken the beam on the REFL OSA was a direct reflection from PRM with the rest of the suspended mirrors misaligned.

One of the upper and lower 11 MHz sidebands is resolved (it is shown at 0.12 sec in the plot) while the other one is still covered by the carrier tail.

The 55 MHz upper and lower sidebands are well resolved (they are at 0.06 and 0.2 sec in the plot).

One of the oscilloscopes monitoring the OSA signals in the control room has a USB interface so that we can record the data into a USB flash memory and plot it like this.

OSArefl.png

Quote from #6375

 I swap an OSA at PSL and OSA at REFL. It was because the PSL-OSA had a better resolution, so we place this better one at REFL. The ND filter (ND3) which was on the way to REFL OSA was replaced by two BSs, because it was producing dirty multiple spots after transmitting.

 

  6377   Wed Mar 7 18:00:39 2012 Ryan FisherSummaryComputer Scripts / ProgramsAlterations to base epics install for installing aLIGO conlog:

Quote:

In order to install the necessary extensions to epics to make the aLIGO conlog work, I have edited one file in the base epics install that affects makefiles:

/cvs/cds/caltech/apps/linux64/epics/base/configure/CONFIG_COMMON

Jamie said he prefers diffs, so I regenerated the original file and did a diff against the current file:

megatron:configure>diff CONFIG_COMMON.orig.reconstructedMar72012 CONFIG_COMMON.bck.Mar72012
206,207c206,210
< USR_CPPFLAGS =
< USR_DBDFLAGS =
---
> USR_CPPFLAGS = -I $(EPICS_BASE)/include
> USR_CPPFLAGS += -I $(EPICS_BASE)/include/os/Linux/
> USR_CPPFLAGS += -I $(EPICS_BASE)/../modules/archive/lib/linux-x86_64/
> USR_DBDFLAGS = -I $(EPICS_BASE)/dbd
> USR_DBDFLAGS += -I $(EPICS_BASE)/../modules/archive/dbd

This is saved in CONFIG_COMMON.diff.Mar72012_1

 After following up with Patrick Thomas for a while trying to make the extensions to epics work within the currently installed epics, he decided that we should just start over with a fresh install of epics 3.14.10.

I am installing this in /ligo/apps/linux-x86_64/epics/base-3.14.10/

Prior to all of this, I had done a lot of installation and configuration of the packages needed to make LAMP work:

sudo apt-get install lamp-server^

sudo apt-get install phpmyadmin

I then continued to follow the instructions on Patrick's wiki:

https://awiki.ligo-wa.caltech.edu/aLIGO/Conlog#EDCU_library

I installed the c_string library into /ligo/apps/linux-x86_64/ according to his instructions.  (prior to my installs, there was no /ligo/ on this machine at all, so I made the needed parent directories with the correct permissions).

That should be everything up to installing epics (working on that now).

  6376   Wed Mar 7 17:39:40 2012 keikoUpdateLSCMICH noise budget on 5 Mar

 This is the calibrated MICH noise budget on Mar 5. There was a sharp peak at 1Hz and a blob on 3 Hz. The demod phase was adjusted for AS55Q.

Mar5-MICHbudget.png

 

Attachment 1: Mar5-MICHbudget.png
Mar5-MICHbudget.png
ELOG V3.1.3-